OpenStudy (anonymous):

Subtract, simplify the result if possible. x 6 x^2-16 x^2+5x+4

OpenStudy (anonymous):

I'm sorry, but what is this asking?

OpenStudy (anonymous):

Oh! I get it. They're fractions.

OpenStudy (anonymous):

hahaha yeah, i don't know how to type it on here. i have to subtract them

OpenStudy (anonymous):

Factorize the denominators x^2-16=(x+4)(x-4) x^2+5x+4=(x+4)(x+1) The rest should be simple :)